# Outlier Gradient Analysis: Efficiently Identifying Detrimental Training Samples for Deep Learning Models

### v3

@@ 1 Introduction
The essence of data-centric learning lies in estimating data influence, also known as data valuation in the context of a learning task (Hammoudeh and Lowd, 2022),. Intuitively, the impact of an individual data sample can be measured by assessing the change in learning utility when training with and without that specific sample. This leave-one-out influence (Cook and Weisberg, 1982) provides a rough gauge of the relative data influence of the specific sample on the otherwise full fixed training set. Shapley value (Ghorbani and Zou, 2019; Jia et al., 2019), originating from cooperative game theory, quantifies the increase in value when a group of samples collaborates to achieve the learning goal. Unlike leave-one-out influence, Shapley value represents the weighted average utility change resulting from adding the sample to different training subsets. Despite the absence of assumptions on the learning model, the aforementioned retraining-based methods incur significant computational costs, especially for large-scale data analysis and deep models (Schioppa et al., 2022).
A popular choice for data valuation applications, such as identifying training samples detrimental to model performance, are influence functions (Koh and Liang, 2017). Essentially, influence functions assess data influence without requiring model retraining. They measure the effect of changing an infinitesimal weight of training samples based on a utility-evaluating function. While influence functions can be accurate or acceptable proxies for convex and certain shallow models, their applicability to deep models is constrained by the strong convexity assumption and the computational cost linked to calculating the inverse of the Hessian matrix (Basu et al., 2020a).
@@
  • Subsequently, we demonstrate the effectiveness of outlier gradient analysis in trimming mislabeled samples from vision datasets across various noise regimes. Additionally, we explore textual applications by conducting experiments on data selection for fine-tuning deep transformer models and identifying influential data for text generation tasks using fine-tuned Large Language Models..
@@ 2 Related Work
Retraining-Based Influence Estimation. Influence estimation approaches can be generally categorized as either retraining-based or gradient-based (Hammoudeh and Lowd, 2022). Retraining-based methods consist of the classical leave-one-out influence approach (Cook and Weisberg, 1982), which consists of removing one training sample at a time, and retraining the model to measure sample influence via performance change. Other representative methods include Shapley value approaches (Ghorbani and Zou, 2019; Jia et al., 2019; Kwon and Zou, 2022), which are model agnostic, but also computationally untenable for large datasets and deep models due to exponential time complexity. Computationally efficient approaches such as KNN-Shap (Jia et al., 2018) can only employ KNN classifiers and hence are not directly applicable to the deep models.
Gradient-Based Influence Estimation. For models trained using gradient descent, gradient-based influence approaches can be used to approximately estimate influence without requiring retraining. The seminal work in this category is that of Koh and Liang (2017), which utilizes a Taylor-series approximation and LiSSA optimization (Agarwal et al., 2017) to compute sample influences. However, the limiting underlying assumption in the formulation is that the model and loss function are convex, which is not true for deep models. Follow-up works such as representer point (Yeh et al., 2018) and Hydra (Chen et al., 2021) inherit these convexity assumptions and suffer from similar issues of applicability. While influence functions have been used for numerous applications in data-centric learning (Feldman and Zhang, 2020; Chhabra et al., 2024; Richardson et al., 2023), they tend to be too computationally expensive for large models, and cannot run in reasonable time. More recently, efficient influence estimation methods such as DataInf (Kwon et al., 2024), Arnoldi iteration (Schioppa et al., 2022), and Kronecker-factored approximation curvature (Grosse et al., 2023) have been proposed which can be employed for large models. Some approaches simply consider the gradients directly as a measure of influence (Pruthi et al., 2020; Charpiat et al., 2019), followed by some ensemble strategies (Bae et al., 2024; Kim et al., 2024). Recent work has also investigated the role of the Hessian and convexity in influence estimation (Schioppa et al., 2024). In contrast, our work aims to circumvent these issues for detrimental sample identification by operating on the gradient space in a skillful manner. Hence, our work paves the way for an efficient and accurate detrimental sample identification framework and adds to the “influence function toolset” for deep models and large datasets. Finally, recent work has also found that self-influence (influence computed on training samples) can be beneficial in detecting detrimental samples (Bejan et al., 2023; Thakkar et al., 2023). For related works on miscellaneous data-centric learning, please refer to Appendix A.
@@ 3.1 Preliminaries on Influence Functions
Let $T$$=$$\{z_{i}\}_{i=1}^{n}$ be a training set, where $z_{i}$ $=$$(x_{i},y_{i})$ includes the input space feature $x_{i}$ and output space label $y_{i}$. A classifier trained using empirical risk minimization on the empirical loss $\ell$ can be written as: $\hat{\theta}$$=$$\operatorname*{arg\,min}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\ell(z_{i};\theta)$. Influence functions (Cook and Weisberg, 1982; Hampel, 1974; Martin and Yohai, 1986) measure the effect of changing an infinitesimal weight of training samples, based on a function that evaluates model utility. Downweighting a training sample $z_{j}$ by a very small fraction $\epsilon$ leads to a model parameter: $\hat{\theta}(z_{j};-\epsilon)=\operatorname*{arg\,min}_{\theta\in\Theta}\frac{1}{n}(\sum_{i=1}^{n}\ell(z_{i};\theta)$--$\epsilon\ell(z_{j};\theta))$. By evaluating the limit as $\epsilon$ approaches 1, the seminal work of Koh and Liang (2017) provides an estimation for the influence score associated with the removal of $z_{j}$ from the training set in terms of training/validation loss as follows:
@@
~$\tilde{\mathcal{I}}(z_{j})=\begin{cases}0\ \textup{(Detrimental Sample)}&\mathcal{I}(z_{j})<0.\\ 1\ \textup{(Beneficial Sample)}&\mathcal{I}(z_{j})\geq 0.\end{cases}$ (2)
~$\tilde{\mathcal{I}}(z_{j})$ can be regarded as a the discrete version of $\mathcal{I}(z_{j})$. Specifically, a value of 0 for ~$\tilde{\mathcal{I}}(z_{j})$ means that removing the sample $z_{j}$ enhances the model’s utility, and that $z_{j}$ is a detrimental sample.
Remark. While influence functions offer a swift estimation for identifying detrimental training samples without the need for costly model retraining, their practical applications to large models are constrained by two prominent drawbacks. The first limitation lies in the necessity of a strictly convex loss function to guarantee the existence of the inverse of the Hessian matrix. The second challenge pertains to the considerable computational expense associated with calculating the inverse of the Hessian. For the first challenge, several possible solutions have been proposed: (1) a convex surrogate model can be used instead of the non-convex model (Chhabra et al., 2024); (2) a damping term can be added to the Hessian to ensure it is positive definite and invertible (Han et al., 2020); and (3) alternative formulations (Basu et al., 2020b; Alaa and Van Der Schaar, 2020) can be used (e.g. the Gauss Newton Hessian (Grosse et al., 2023) instead of the standard Hessian). Note that some studies bypass the convexity assumption and directly apply influence functions to deep models, yielding effective results. (Grosse et al., 2023). For the second challenge, various matrix inverse techniques are employed to expedite the computation process, including LiSSA optimization (Koh and Liang, 2017) and swapping the order of the matrix inversion (Kwon et al., 2024), among several others. Considerable efforts have been dedicated to addressing the aforementioned challenges with promising results– however, in this paper we target the second challenge for identifying/removing detrimental samples.
@@ 3.3 Our Approach: Outlier Gradient Analysis
1:Input: Training set $T$, loss $\ell$, trained model param $\hat{\theta}$, outlier analysis algorithm $\mathcal{A}$, trimming budget $k$
@@ 4 Hypothesis Verification on Synthetic Data
We also conduct a quantitative evaluation to assess the advantages of our approach compared to three recently proposed noisy label correction methods and six influence function-based approaches, as detailed in Table 1. Specifically, we measure ground-truth outlier predictive accuracy and the performance gain achieved by removing detrimental samples. For noisy label correction approaches we consider: Normalized Margin (Northcutt et al., 2021), Self-Confidence (Müller and Markert, 2019), and Confidence-Weighted Entropy (Kuan and Mueller, 2022). The influence function approaches include computing the Hessian exactly (Cook and Weisberg, 1982), using the Hessian-free gradient tracing approach [^2] We only use the last checkpoint in Gradient Tracing (Pruthi et al., 2020) for fair comparisons. by (Pruthi et al., 2020), LiSSA-based optimization (Koh and Liang, 2017), the recently proposed influence estimation approach DataInf (Kwon et al., 2024), self-influence using LiSSA as in (Bejan et al., 2023), and self-influence using DataInf. We compute influences only using the training samples and performance is measured on the test set.
Our outlier gradient analysis approaches demonstrate high accuracy in identifying mislabeled outliers (96-98%), outperforming all three noisy label correction baselines (only 82% accuracy) and among influence baselines, all exhibit similar performance except for exact Hessian computation, which attains 90% accuracy. Next, we evaluate model performance gain by removing detected outlier samples and retraining the MLP on the trimmed dataset. Here the benefits of our iForest outlier gradient analysis can be observed, as it increases performance from 90% to 96% while the overtly simple L1/L2-norm outlier analysis approaches are not as effective. The other baselines exhibit performance variations between 89-91%. This emphasizes the effectiveness of our iForest approach, while exhibiting low time complexity (refer to Appendix C.3 for details on computational complexity).
@@ 5 Noisy Label Correction for Vision Datasets
Here we demonstrate the effectiveness of our approach in addressing noisy label correction using the CIFAR-10N and CIFAR-100N real-world noisy label datasets (Wei et al., 2022). These datasets stem from the original CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009), but introduce label inaccuracies due to crowdsourced labeling. CIFAR-10N has 3 different noise settings: Aggregate, Random, and Worst– these correspond to majority voting across 3 annotators, the first annotator label, and selecting the worst annotator label, respectively. CIFAR-100N only has a single noise setting.
@@
Figure 2: Some detrimental samples detected using our proposed outlier gradient analysis. Top row: CIFAR-10N; bottom row: CIFAR-100N. Top label (red): noisy label; bottom label (green): correct class.
Additional Analyses. We conduct ablations on the iForest parameters in Appendix C.4. Further, we provide running time experiments on CIFAR-10N and CIFAR-100N in Appendix C.3 along with the other baselines. We also provide results with ResNet-18 as the base model in Appendix C.5 and on ImageNet (Deng et al., 2009) in Appendix C.6, showing similar trends. Finally, approaches for noisy learning can be categorized into methods that either change the loss function or model architecture or methods that identify noisy samples and remove/relabel them for improving performance (Algan and Ulusoy, 2021). Since our approach belongs to the latter category, we only compare against other approaches from this category. For completeness, we also present results comparing our approach with some others in the former category in Appendix C.7. We would like to emphasize that this is not an exhaustive list of baselines and noisy learning by adjusting the loss/model is not the focus of our work (but detecting detrimental samples is). Moreover, our algorithm could also be combined with approaches from both categories for additional gains.
@@ 6 Data Selection for Fine-tuning NLP Models
Figure 3: Performance of the data selection task utilizing outlier gradient trimming and other influence-based baselines for fine-tuning RoBERTa via LoRA over 3 runs.
The results over three runs are presented in Figure 3 for all four GLUE datasets. We only show trends for iForest based outlier gradient analysis to aid visualization since performance is similar for the L1/L2-norm methods. It can be seen that our outlier gradient trimming approach markedly outperforms all other baselines but the Self-LiSSA (Bejan et al., 2023) self-influence baseline is competitive with our approach. More specifically, outlier gradient analysis achieves slightly better test set results on QNLI, SST2, QQP, and on MRPC, Self-LiSSA and outlier gradient analysis are on par with each other. Here, we would like to emphasize that despite competitive performance, our outlier gradient analysis is orders of magnitude faster than Self-LiSSA, as shown in experiments of Appendix C.3. These results highlight the effectiveness of our proposed outlier gradient analysis/trimming approach in selecting relevant data for fine-tuning NLP models while being more computationally efficient.
@@ 7 Extending to Influential Data Identification for LLMs
We now consider an alternate task– demonstrating the effectiveness of our proposed outlier gradient analysis in identifying influential data samples for Large Language Models (LLMs), using the proposed benchmarks from DataInf (Kwon et al., 2024). The LLM influential data identification task at its core is a similarity measurement task, as it seeks to ascertain which fine-tuning prompts are most similar to a given test sample. More specifically, the goal is to assess what training set prompts (used for LoRA fine-tuning) are most influential for a given unseen test prompt. The robustness and effectiveness of influence estimation are gauged based on whether the identified training set prompts belong to the same class category as the given test prompt. We utilize the three benchmark datasets introduced in DataInf (Kwon et al., 2024): Sentence Transformations, Math Without Reasoning, and Math With Reasoning, to conduct the influential data identification experiment on the Llama-2-13B-chat [^3] https://ai.meta.com/llama/. LLM. For each of the influence identification benchmark datasets, there are 900 training samples for LoRA fine-tuning, and 10 categories or classes of task types with 90 samples belonging to each class. For each dataset there are 100 test set prompts with 10 test set prompts per class category.
Table 3: AUC/Recall for outlier gradient analysis and baselines for influential class detection on three text generation tasks on Llama2-13B-chat LLM.
@@
In Kwon et al. (2024), to predict the most influential training samples given a test set prompt, the authors assign a pseudo label to every data point in the training set (1 if it is in the same class/task category as the test data prompt, or 0 otherwise). This set serves as a ground-truth for measuring performance of identifying influential data samples. Next, they calculate the Area Under the Curve (AUC) by comparing the absolute values of the influence function (for each training set prompt corresponding to a given test prompt) with these pseudo labels. Clearly, a high AUC signifies that training data samples from the same category have a significant influence on the given test prompt. The average AUC across all test data points is then recorded, and is denoted as the Class Detection (AUC) metric. Additionally, another metric is used– for every test data prompt, the authors determine if the proportion of training data prompts belonging to the same class/category are within the top 90 (# of training prompts in each category) influential samples. The average % across all test data points is calculated and this metric is denoted as Class Detection (Recall), where higher recall is better.
As part of this task, we need to measure similarity between train and test set samples. Note that for our experiments on identifying detrimental samples outlier gradient analysis only operated on the training set (i.e., it uses the training set gradients). However, to extend outlier analysis to this task while maintaining consistency with the previous experiments and methods, we will train 10 individual iForest estimators for each class prompt category, as the ultimate objective is to use outlier gradient analysis for prompt class detection. Each class’s iForest estimator is trained solely on the gradient space of training prompts from that category. Subsequently, for each test set prompt, we utilize each iForest estimator to generate an outlier score based on the gradient space of that test sample. This enables us to conduct the influential data identification experiment for our proposed method. The other baseline influence methods already have access to the given test set sample and can use that information directly for analyzing which training sample is most influential.

### v5

@@ 1 Introduction
The essence of data-centric learning lies in estimating data influence, also known as data valuation in the context of a learning task (Hammoudeh & Lowd, 2022). Intuitively, the impact of an individual data sample can be measured by assessing the change in learning utility when training with and without that specific sample. This leave-one-out influence (Cook & Weisberg, 1982) provides a rough gauge of the relative data influence of the specific sample on the otherwise full fixed training set. Shapley value (Ghorbani & Zou, 2019; Jia et al., 2019), originating from cooperative game theory, quantifies the increase in value when a group of samples collaborates to achieve the learning goal. Unlike leave-one-out influence, Shapley value represents the weighted average utility change resulting from adding the sample to different training subsets. Despite the absence of assumptions on the learning model, the aforementioned retraining-based methods incur significant computational costs, especially for large-scale data analysis and deep models (Schioppa et al., 2022).
A popular choice for data valuation applications, such as identifying training samples detrimental to model performance, are influence functions (Koh & Liang, 2017). Essentially, influence functions assess data influence without requiring model retraining. They measure the effect of changing an infinitesimal weight of training samples based on a utility-evaluating function. While influence functions can be accurate or acceptable proxies for convex and certain shallow models, their applicability to deep models is constrained by the strong convexity assumption and the computational cost linked to calculating the inverse of the Hessian matrix (Basu et al., 2020a).
@@
  • Subsequently, we demonstrate the effectiveness of outlier gradient analysis in trimming mislabeled samples from vision datasets across various noise regimes. Additionally, we explore textual applications on data selection for fine-tuning deep transformer models and identifying influential data for text generation tasks using fine-tuned Large Language Models.
@@ 2 Related Work
Retraining-Based Influence Estimation. Influence estimation approaches can be generally categorized as either retraining-based or gradient-based (Hammoudeh & Lowd, 2022). Retraining-based methods consist of the classical leave-one-out influence approach (Cook & Weisberg, 1982), which consists of removing one training sample at a time, and retraining the model to measure sample influence via performance change. Other representative methods include Shapley value approaches (Ghorbani & Zou, 2019; Jia et al., 2019; Kwon & Zou, 2022), which are model agnostic, but also computationally untenable for large datasets and deep models due to exponential time complexity. Computationally efficient approaches such as KNN-Shap (Jia et al., 2018) can only employ KNN classifiers and hence are not directly applicable to the deep models.
Gradient-Based Influence Estimation. For models trained using gradient descent, gradient-based influence approaches can be used to approximately estimate influence without requiring retraining. The seminal work in this category is that of (Koh & Liang, 2017), which utilizes a Taylor-series approximation and LiSSA optimization (Agarwal et al., 2017) to compute sample influences. However, the limiting underlying assumption in the formulation is that the model and loss function are convex, which is not true for deep models. Follow-up works such as representer point (Yeh et al., 2018) and Hydra (Chen et al., 2021) inherit these convexity assumptions and suffer from similar issues of applicability. While influence functions have been used for numerous applications in data-centric learning (Feldman & Zhang, 2020; Chhabra et al., 2024; Richardson et al., 2023), they tend to be too computationally expensive for large models, and cannot run in reasonable time. More recently, efficient influence estimation methods such as DataInf (Kwon et al., 2024), Arnoldi iteration (Schioppa et al., 2022), and Kronecker-factored approximation curvature (Grosse et al., 2023) have been proposed which can be employed for large models. Some approaches simply consider the gradients directly as a measure of influence (Pruthi et al., 2020; Charpiat et al., 2019), followed by some ensemble strategies (Bae et al., 2024; Kim et al., 2024). Recent work has also investigated the role of the Hessian and convexity in influence estimation (Schioppa et al., 2024). In contrast, our work aims to circumvent these issues for detrimental sample identification by operating on the gradient space in a skillful manner. Hence, our work paves the way for an efficient and accurate detrimental sample identification framework and adds to the “influence function toolset” for deep models and large datasets. Finally, recent work has also found that self-influence (influence computed on training samples) can be beneficial in detecting detrimental samples (Bejan et al., 2023; Thakkar et al., 2023). For related works on miscellaneous data-centric learning, please refer to Appendix A.
@@ 3.1 Preliminaries on Influence Functions
Let $T$$=$$\{z_{i}\}_{i=1}^{n}$ be a training set, where $z_{i}$ $=$$(x_{i},y_{i})$ includes the input space feature $x_{i}$ and output space label $y_{i}$. A classifier trained using empirical risk minimization on the empirical loss $\ell$ can be written as: $\hat{\theta}$$=$$\operatorname*{arg\,min}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\ell(z_{i};\theta)$. Influence functions (Cook & Weisberg, 1982; Hampel, 1974; Martin & Yohai, 1986) measure the effect of changing an infinitesimal weight of training samples, based on a function that evaluates model utility. Downweighting a training sample $z_{j}$ by a very small fraction $\epsilon$ leads to a model parameter: $\hat{\theta}(z_{j};-\epsilon)=\operatorname*{arg\,min}_{\theta\in\Theta}\frac{1}{n}(\sum_{i=1}^{n}\ell(z_{i};\theta)$--$\epsilon\ell(z_{j};\theta))$. By evaluating the limit as $\epsilon$ approaches 1, the seminal work of Koh & Liang (2017) provides an estimation for the influence score associated with the removal of $z_{j}$ from the training set in terms of training/validation loss as follows:
@@
~$\tilde{\mathcal{I}}(z_{j})=\begin{cases}0\ \textup{(Detrimental Sample)}&\mathcal{I}(z_{j})<0.\\ 1\ \textup{(Beneficial Sample)}&\mathcal{I}(z_{j})\geq 0.\end{cases}\vspace{-2mm}$ (2)
~$\tilde{\mathcal{I}}(z_{j})$ can be regarded as the discrete version of $\mathcal{I}(z_{j})$. Specifically, a value of 0 for ~$\tilde{\mathcal{I}}(z_{j})$ means that removing the sample $z_{j}$ enhances the model’s utility, and that $z_{j}$ is a detrimental sample.
Remark. While influence functions offer a swift estimation for identifying detrimental training samples without the need for costly model retraining, their practical applications to large models are constrained by two prominent drawbacks. The first limitation lies in the necessity of a strictly convex loss function to guarantee the existence of the inverse of the Hessian matrix. The second challenge pertains to the considerable computational expense associated with calculating the inverse of the Hessian. For the first challenge, several possible solutions have been proposed: (1) a convex surrogate model can be used instead of the non-convex model (Chhabra et al., 2024); (2) a damping term can be added to the Hessian to ensure it is positive definite and invertible (Han et al., 2020); and (3) alternative formulations (Basu et al., 2020b; Alaa & Van Der Schaar, 2020) can be used (e.g. the Gauss Newton Hessian (Grosse et al., 2023) instead of the standard Hessian). Note that some studies bypass the convexity assumption and directly apply influence functions to deep models, yielding effective results. (Grosse et al., 2023). For the second challenge, various matrix inverse techniques are employed to expedite the computation process, including LiSSA optimization (Koh & Liang, 2017) and swapping the order of the matrix inversion (Kwon et al., 2024), among several others. Considerable efforts have been dedicated to addressing the aforementioned challenges with promising results– however, in this paper we target the second challenge for identifying/removing detrimental samples.
@@ 3.3 Our Approach: Outlier Gradient Analysis
1:  Input: Training set $T$, loss function $\ell$, model parameter $\hat{\theta}$, outlier analysis algorithm $\mathcal{A}$, trimming budget $k$
@@ 4 Hypothesis Verification on Synthetic Data
We also conduct a quantitative evaluation to assess the advantages of our approach compared to three recently proposed noisy label correction methods and six influence function-based approaches, as detailed in Table 1. Specifically, we measure ground-truth outlier predictive accuracy and the performance gain achieved by removing detrimental samples. For noisy label correction approaches we consider: Normalized Margin (Northcutt et al., 2021), Self-Confidence (Müller & Markert, 2019), and Confidence-Weighted Entropy (Kuan & Mueller, 2022). The influence function approaches include computing the Hessian exactly (Cook & Weisberg, 1982), using the Hessian-free gradient tracing approach by (Pruthi et al., 2020), LiSSA-based optimization (Koh & Liang, 2017), the recently proposed influence estimation approach DataInf (Kwon et al., 2024), self-influence using LiSSA as in Bejan et al. (2023), and self-influence using DataInf. We compute influences only using the training samples and performance is measured on the test set.
Our outlier gradient analysis approaches demonstrate high accuracy in identifying mislabeled outliers (96-98%), outperforming all three noisy label correction baselines (only 82% accuracy) and among influence baselines, all exhibit similar performance except for exact Hessian computation, which attains 90% accuracy. Next, we evaluate model performance gain by removing detected outlier samples and retraining the MLP on the trimmed dataset. Here the benefits of our iForest outlier gradient analysis can be observed, as it increases performance from 90% to 96% while the overtly simple L1/L2-norm outlier analysis approaches are not as effective. The other baselines attain performance between 89-91%. This emphasizes the effectiveness of our iForest approach, while exhibiting low time complexity (refer to Appendix C.3 for details on computational complexity).
@@ 5 Noisy Label Correction for Vision Datasets
We now demonstrate the effectiveness of our approach in addressing noisy label correction using the CIFAR-10N and CIFAR-100N real-world noisy label datasets (Wei et al., 2022). These datasets stem from the original CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009), but introduce label inaccuracies due to crowdsourced labeling. CIFAR-10N has 3 different noise settings: Aggregate, Random, and Worst– these correspond to using majority voting across 3 annotators, first annotator label, and worst annotator label, respectively. CIFAR-100N only has one noise setting.
@@
Figure 2: Detrimental samples detected using our outlier gradient analysis. Top row: CIFAR-10N; bottom row: CIFAR-100N. Top label (red): noisy label; bottom label (green): correct class.
Additional Analyses. We conduct ablations on the iForest parameters in Appendix C.4. Further, we provide running time experiments on CIFAR-10N and CIFAR-100N in Appendix C.3 along with the other baselines. We also provide results with ResNet-18 as the base model in Appendix C.5 and on ImageNet (Deng et al., 2009) in Appendix C.6, showing similar trends. Finally, approaches for noisy learning can be categorized into methods that either change the loss function or model architecture or methods that identify noisy samples and remove/relabel them for improving performance (Algan & Ulusoy, 2021). Since our approach belongs to the latter category, we only compare against other approaches from this category. For completeness, we also present results comparing our approach with some others in the former category in Appendix C.7. We would like to emphasize that this is not an exhaustive list of baselines and noisy learning by adjusting the loss/model is not the focus of our work (but detecting detrimental samples is). Moreover, our algorithm could also be combined with approaches from both categories for additional gains. Finally, we also conducted experiments using two new influence function methods: TRAK (Park et al., 2023) and GEX (Kim et al., 2024). While we were able to obtain results for CIFAR-10N (please refer to Appendix C.8 for results), both methods got out-of-memory errors on CIFAR-100N for the same experimental set-up as other influence methods. Given their shortcomings, we did not consider them for the other experiments.
@@ 6 Data Selection for Fine-tuning NLP Models
Figure 3: Performance of the data selection task using outlier gradient trimming and other influence baselines for fine-tuning RoBERTa.
@@
The results over three runs are presented in Figure 3 for all four GLUE datasets. We only show trends for iForest based outlier gradient analysis to aid visualization since performance is similar for the L1/L2-norm methods. It can be seen that our outlier gradient trimming approach markedly outperforms all other baselines– more specifically, outlier gradient analysis achieves the best test set results on QNLI, SST2, QQP, and on MRPC, Self-LiSSA (Bejan et al., 2023) and outlier gradient analysis are on par with each other. Despite this competitive performance, our outlier gradient analysis is orders of magnitude faster than Self-LiSSA, as shown in experiments of Appendix C.3. This highlights the effectiveness of our proposed outlier gradient analysis approach in selecting relevant data for fine-tuning NLP models while being more computationally efficient.
@@ 7 Extending to Influential Data Identification for LLMs
We now consider an alternate task– demonstrating the effectiveness of our proposed outlier gradient analysis in identifying influential data samples for Large Language Models (LLMs), using the proposed benchmarks from DataInf (Kwon et al., 2024). The LLM influential data identification task at its core is a similarity measurement task, as it seeks to ascertain which fine-tuning prompts are most similar to a given test sample (Askari et al., 2025). More specifically, the goal is to assess what training set prompts (used for LoRA fine-tuning) are most influential for a given unseen test prompt. The robustness and effectiveness of influence estimation are gauged based on whether the identified training set prompts belong to the same class category as the given test prompt. We utilize the three benchmark datasets introduced in DataInf (Kwon et al., 2024): Sentence Transformations, Math Without Reasoning, and Math With Reasoning, to conduct the influential data identification experiment on the Llama-2-13B-chat LLM (Touvron et al., 2023). For each of the influence identification benchmark datasets, there are 900 training samples for LoRA fine-tuning, and 10 categories or classes of task types with 90 samples belonging to each class. For each dataset there are 100 test set prompts with 10 test set prompts per class category.
In (Kwon et al., 2024), to predict the most influential training samples given a test set prompt, the authors assign a pseudo label to every data point in the training set (1 if it is in the same class/task category as the test data prompt, or 0 otherwise). This set serves as a ground-truth for measuring performance of identifying influential data samples. Next, they calculate the Area Under the Curve (AUC) by comparing the absolute values of the influence function (for each training set prompt corresponding to a given test prompt) with these pseudo labels. Clearly, a high AUC signifies that training data samples from the same category have a significant influence on the given test prompt. The average AUC across all test data points is then recorded, and is denoted as the Class Detection (AUC) metric. Additionally, another metric is used– for every test data prompt, the authors determine if the proportion of training data prompts belonging to the same class/category are within the top 90 (# of training prompts in each category) influential samples. The average % across all test data points is calculated and this metric is denoted as Class Detection (Recall), where higher recall is better.
Table 3: AUC/Recall for outlier gradient analysis and baselines for influential class detection for three tasks on Llama2-13B LLM.
@@
As part of this task, we need to measure similarity between train and test set samples. Note that for our experiments on identifying detrimental samples outlier gradient analysis only operated on the training set (i.e., it uses the training set gradients). However, to extend outlier analysis to this task while maintaining consistency with the previous experiments and methods, we will train 10 individual iForest estimators for each class prompt category, as the ultimate objective is to use outlier gradient analysis for prompt class detection. Each class’s iForest estimator is trained solely on the gradient space of training prompts from that category. Subsequently, for each test set prompt, we utilize each iForest estimator to generate an outlier score based on the gradient space of that test sample, enabling us to conduct the influential data identification experiment. Note that the other baseline influence methods already have access to the given test set sample and can use that information directly for analyzing which training sample is most influential.